OpenStudy (clamin):

Each equation given below describes a parabola. Which statement best compares their graphs? x= 5y^2 x=3y^2 a. both parabolas open upward and x=5y^2 is wider than x= 3y^2 b. both parabolas open to the righ and x=3y^2 is wider than x=5y^2 . c. both parabolas open upward and x=3y^2 is wider than x=5y^2 d. both parabolas open to the right and x=5y^2 is wider than x=3y^2

OpenStudy (tkhunny):

Well, what say you? You MUST have SOME opinion.

OpenStudy (clamin):

i cant choose between a and c

OpenStudy (tkhunny):

That is good, since neither of those is correct. If they were y = x^2, we should be talking about A or C. Since these are x = y^2, we should be talking about B or D.

OpenStudy (clamin):

d @tkhunny

OpenStudy (tkhunny):

You should learn to speak in sentences. 3 is wider than 5.

OpenStudy (danjs):

think about it for a minute x = 3*y^2 , x = 5*y^2 when you decide a y value here, say y=3, the x coordinate will change depending on that constant factor, the 5 makes x move larger faster, overall the larger 5 will create a "skinnier" parabola

OpenStudy (danjs):

as x grows larger at a quicker rate

OpenStudy (welshfella):